A coffee machine dispenses coffee according to a
normal distribution, with a mean of 8.4 ounces and
a standard deviation of 0.16 ounces. Find the
probabilities of the following events.
a. A cup is dispensed with less than 8 ounces.
b. A cup has at least 8.5 ounces.

Answer by rothauserc(4718)

a. z-score = (8 - 8.4) / 0.16 = −2.5
lookup z-score in normal distribution z-tables
Probability ( X < 8 ) = 0.0062
:
b. Probability ( X > or = 8.5 ) = 1 - Probability ( X < 8.5 )
z-score = (8.5 - 8.4) / 0.16 = 0.625
Probability ( X < 8.5 ) = 0.73401
Probability ( X > or = 8.5 ) = 1 - 0.73401 = 0.266
: